is (5,12)a solution to the equation: y = − 4x + 3. Prove your answer!
Found 2 solutions by fractalier, Theo:
Answer by fractalier(6550)

We substitute in for x and y and see if we have a true statement...like this...
y = -4x + 3
12 = -4(5) + 3
12 = -20 + 3
12 = -17
12 does not equal -17 so it is NOT a solution.

Answer by Theo(13342)

(5,12) means x = 5 when y = 12, or y = 12 when x = 5.

equation is y = -4x + 3

when x = 5, y = -4x + 3 becomes y = -4*5 + 3 which becomes y = -20 + 3 which becomes y = -17.

when y = 12, y = -4x + 3 becomes 12 = -4x + 3.
subtract 3 from both sides of this equation to get 9 = -4x
divide both sides of this equation by -4 to get x = -9/4.

(5,12) is not a solution to the equation of y = -4x + 3.